OpenStudy (anonymous):

find two consecutive even integers such that the larger, added to eight times the smaller, equals 110.

OpenStudy (saifoo.khan):

x+2+8=110 YOU FORGOT THAT THE Y MULTIPLIER FOR THE 8. LET x & x=2 BE THE 2 INTEGERS. (x+2)+8x=110 9x+2=110 9x=110-2 9x=108 x=108/9 x=12 ANS. PROOF: 12+2=14 ANS. FOR THE LARGER INTEGER.

OpenStudy (anonymous):

wooow someone looks smart